# Depth-Based Local Center Clustering: A Framework for Handling Different Clustering Scenarios

### v1

@@ Abstract
Cluster analysis, or clustering, plays a crucial role across numerous scientific and engineering domains. Despite the wealth of clustering methods proposed over the past decades, each method is typically designed for specific scenarios and presents certain limitations in practical applications. In this paper, we propose depth-based local center clustering (DLCC). This novel method makes use of data depth, which is known to produce a center-outward ordering of sample points in a multivariate space. However, data depth typically fails to capture the multimodal characteristics of data, something of the utmost importance in the context of clustering. To overcome this, DLCC makes use of a local version of data depth that is based on subsets of data. From this, local centers can be identified as well as clusters of varying shapes. Furthermore, we propose a new internal metric based on density-based clustering to evaluate clustering performance on non-convex clusters. Overall, DLCC is a flexible clustering approach that seems to overcome some limitations of traditional clustering methods, thereby enhancing data analysis capabilities across a wide range of application scenarios.
Index Terms:
Clustering, statistical depth, similarity matrix, Mahalanobis depth, spatial depth, internal metric.
@@ I Introduction
The complexity of data distributions has led to the development of numerous clustering algorithms, each having a unique set of strengths and constraints. We next briefly discuss different traditional clustering methods to highlight their distinct characteristics and limitations. In this, and in what follows, we consider the objects to be clustered as points in the $d$ dimensional real space $\mathbb{R}^{d}$.
@@ I-A1 Center-Based Clustering
Center-based clustering methods, with k-means clustering [9] being a prime example, are often utilized due to their simplicity and effectiveness. However, k-means clustering has several limitations. Firstly, it is sensitive to outliers as it uses the means of the points in each cluster as cluster centers. Secondly, even in outlier-free contexts, this approach is most adequate when clusters are spherically symmetric, which may not always be representative of the actual data distribution. Lastly, the algorithm is sensitive to initialization and can easily get trapped in local minima. To address some of these issues, several variants of the k-means algorithm have been developed:
  • Partitioning around medoids (PAM; [10]), was introduced to combat the sensitivity to outliers. Instead of cluster means, PAM uses medoids (a form of multivariate median) as cluster centres.
@@
  • Weighted k-means (W-k-means; [11]) and lasso weighted k-means (LW-k-means; [12]) were designed to address the underlying spherical symmetry assumption for clusters. These both provide a more flexible k-means approach by incorporating a weight vector for each variable. LW-k-means improves W-k-means by adding a lasso penalty to the objective function, enabling feature selection in high-dimensional data.
  • Power k-means [13] and kernel power k-means [14] were developed to address the sensitivity of k-means to initialization and tendency to get trapped in local minima by redefining the objective function using power means. Kernel power k-means extends this by incorporating multiple kernel configurations to handle non-linear separable data via higher-dimensional projections.
I-A2 Density-Based and Hierachical Clustering
@@
Density-based clustering assumes that clusters are regions of high density separated by low-density areas, making it well-suited for detecting arbitrarily shaped clusters. For instance, density-based spatial clustering of applications with noise (DBSCAN; [15]) sets a threshold and minimum point count for a region to be considered dense. Then, dense regions that can be directly connected with each other are merged into a cluster. However, DBSCAN’s reliance on a fixed threshold limits its ability to handle clusters with varying densities.
Hierarchical clustering, on the other hand, adopts a tree-like approach to iteratively merge or split clusters. It offers two main strategies: agglomerative (bottom-up) and divisive (top-down). The resulting dendrogram represents cluster relationships, from an all-inclusive cluster at the top to individual objects at the bottom [16]. While hierarchical clustering provides an intuitive structure, its high computational cost and irreversible decisions make it less suitable for large datasets or complex cluster shapes.
Density peak clustering (DPC) [17], introduced in $2014$, takes a different approach by combining density and distance to identify cluster centers. It locates density peaks based on a decision graph, where density is plotted on the y-axis and the distance to the nearest higher-density point on the x-axis. This method allows for manual selection of cluster centers and performs well with varying-density clusters. However, it can struggle with robustness, as density peaks may correspond to satellite points rather than true cluster centers.
An improved version, fast main density peak clustering within relevant regions (R-MDPC) [18], proposed in $2024$, incorporates hierarchical concepts by applying single-linkage clustering to density peaks, defining relevant regions as density-connected areas. This enhancement better handles complex cluster shapes by merging single-density peak clusters and reducing satellite peak interference. However, for highly overlapping or extremely high-dimensional datasets, R-MDPC may collapse all points into one region, behaving similarly to the original DPC.
I-A3 Model-Based Clustering
Model-based clustering typically assumes the points to be clustered stem from an underlying mixture distribution [19], whose components are naturally associated with the clusters to be constructed. For example, working from a Gaussian mixture model (GMM), one can identify elliptical clusters with varying orientations, shapes, and sizes. Other mixture models handling asymmetric distributions have also been developed, see, e.g., [20], [21], [22], and [23]. These different families of mixture models lead to clustering methods typically relying on the expectation-maximization (EM) algorithm [24] to find optimal parameter values, which can be computationally intensive. Moreover, selection of the most appropriate family of mixture models can be challenging unless knowledge about the data-producing mechanism and/or subject area expertise is available.
I-A4 Spectral and Subspace Clustering
Spectral clustering [25], a notable graph-based method, treats clustering as a graph partitioning problem, converting data points into nodes and their similarities into connecting edge weights. It is well-suited for high-dimensional datasets, often utilizing a sparse adjacency matrix derived from the k-nearest neighbors methods. After representing the data as a similarity graph, the clustering problem can essentially be seen as partitioning the graph in such a way that edges between groups are given low weights and edges within groups high weights [26].
Spectral clustering has been extended and has led to other clustering techniques. It is, for instance, an essential component of modern subspace clustering algorithms. These algorithms have gained attention due to their capacity for identifying low-dimensional representations of high-dimensional data [27]. In particular, sparse subspace clustering (SSC) [28] is a key evolution in this domain. Making use of the self-expressiveness property, points are expressed as linear combinations of other points within the same subspace. Therefore, SSC merely requires the coefficient matrix of these linear combinations. As a result, SSC overcomes two major challenges of spectral clustering: selecting an appropriate neighborood size and handling points near the intersection of two subspaces. However, for low-dimensional data, the premise that data points can be effectively represented as linear combinations of the lower-dimensional subspace may falter, potentially leading to a diminished efficacy of SSC. This makes SSC a tool that is primarily designed for high-dimensional datasets.
@@ I-B Motivation
Our objective is to develop a versatile clustering framework capable of addressing diverse challenges, from convex to non-convex cluster shapes, and from low- to high-dimensional data. A common theme across clustering methods is their reliance on representative points, or exemplars. Center-based methods define exemplars by minimizing within-cluster distances, often relying on distance metrics. However, a specific distance measure often performs well only for certain types of data, and adapting to more diverse structures may require mapping data into higher-dimensional spaces via kernel methods. The effectiveness of such approaches, however, hinges on selecting an appropriate kernel, which can be challenging and highly sensitive to the data. Density-based methods, on the other hand, define exemplars through local density variations. These methods identify clusters by connecting dense regions, but the appropriate definition of density often depends on the dimensionality of the data. For example, density measures that work well in low-dimensional spaces may lose effectiveness in high-dimensional settings due to the sparsity of data. Moreover, in overlapping regions, density connections often struggle to distinguish between clusters. Both center-based and density-based approaches share a reliance on exemplars but are inherently limited by their respective definitions of exemplars.
A versatile clustering framework must overcome these limitations by introducing an adaptive similarity measure that avoids strong parametric assumptions and generalizes across diverse data geometries. Equally important is the need for a robust exemplar definition that is consistent across dimensions and cluster shapes, along with effective grouping strategies for these exemplars to form meaningful clusters.
@@
In response, our method introduces a more flexible exemplar definition by leveraging statistical data depth, a nonparametric tool for multivariate analysis. Tukey [29] pioneered the idea of center-outward ordering in multivariate data, with the introduction of the first depth definition, half-space depth. Data depth enables the center-outward ordering of data, making it a potent tool for locating a center for any dataset. When considering subsets of the data, data depth with respect to each subset aids in defining multiple exemplars. Naturally, as the subset size grows, the number of exemplars diminishes, and vice versa.
@@
Building on this concept, we propose the depth-based local center clustering (DLCC) algorithm. Equipped with two grouping strategies for depth-defined representative points, DLCC inherits strengths from both center-based and density-based methods, while offering greater flexibility to tackle a wide range of clustering problems. Here, the term “depth-based” encompasses two aspects: (1) data depth is used to construct a similarity matrix instead of relying on distance, adaptively capturing the geometric characteristics of each point’s location; (2) data depth measures the centrality of points within each point’s neighborhood, helping to identify exemplars.
Furthermore, ground truth is typically not available in real-world clustering applications, necessitating the use of internal metrics to evaluate performance. Established metrics like the average silhouette width (ASW; [49]), which measures the average distance within clusters relative to the nearest cluster, and the Calinski-Harabasz index (CHI; [50]), which relies on distances from points to their cluster centroids, inherently favor spherical clusters. This preference limits their effectiveness in capturing non-convex shapes, as illustrated in Figure 1, where they will incorrectly favor clustering result (c) over (a). To address these shortcomings, we propose a new internal metric, the depth-based clustering (DC) metric, designed to more accurately handle complex cluster shapes and assist in parameter selection for our algorithm. Further details will be discussed in Section IV-B.
Figure 1: Clustering results for the synthetic dataset SmileyF. Panel (a) shows the ground truth, (b) misclusters several points in the eyes, and (c) fails to capture the face pattern.
I-C Main Contributions and Organization of the Paper
@@
  • Depth-Based Similarity Matrix: We improve the efficiency when constructing a depth-based similarity matrix using Mahalanobis depth (MD) and Spatial depth (SD). For the MD-based similarity matrix, model-based clustering methods are employed to predefine covariance matrices that approximate the data’s distribution shapes. For the SD-based similarity matrix, we develop an efficient matrix-based computation method.
  • DLCC Clustering Framework: We develop a novel clustering framework, DLCC, which includes the definition of a similarity measure and local centers, and the introduction of methods for filtering and grouping local centers. Experiments shows it achieves competitive performance across different types of clustering challenges among the representative algorithms from each type of clustering method referenced in Section I-A.
  • Internal Evaluation Metric: We propose a depth-based clustering metric as an internal evaluation criterion. This metric combines the density-based clustering concept with the depth-based similarity matrix to evaluate clustering performance, particularly for non-convex cluster structures.
The remainder of the paper is organized as follows: Section II reviews key concepts of data depth and introduces the depth-based similarity matrix. Section III provides a detailed description of the DLCC framework. In Section IV, we discuss the hierarchical structure of DLCC and present the proposed internal metric for parameter selection. Section V showcases experimental results of DLCC using both synthetic and real-world datasets. Lastly, Section VI concludes the paper.
@@
DatasetDepth-basedsimilarity matrixSection IINeighborhoods &Local centersSection III Filtered centersSection III-AGroups of filtered centersSection III-A &\&& Section IVTemporaryclustersSection III-BUnlabelled pointsSection III-BFinal clustersclassification
@@
$D_{\mathrm{MD/SD}}(\mathbf{z}|\mathbf{X})$ Mahalanobis or spatial depth of $\mathbf{z}$ with respect to $\mathbf{X}$;
@@
$\mathbb{S}$ Depth-based similarity matrix, with $\mathbb{S}_{ij}$ its $ij$-th entry;
@@
$c$, $\mathbf{c}$, $\mathbf{C}$ A local center, a certain set of local centers, and the set of all filtered centers, respectively;
$\mathbb{M}$ Local center neighborhoods’ similarity matrix, with $\mathbb{M}_{ij}$ its $ij$-th entry;
@@
$G$ and $K$ Number of groups and clusters, typically $G=K$;
$k$ Cluster index;

### v2

@@ Abstract
Cluster analysis, or clustering, plays a crucial role across numerous scientific and engineering domains. Despite the wealth of clustering methods proposed over the past decades, each method is typically designed for specific scenarios and presents certain limitations in practical applications. In this paper, we propose depth-based local center clustering (DLCC). This novel method makes use of data depth, which is known to produce a center-outward ordering of sample points in a multivariate space. However, data depth typically fails to capture the multimodal characteristics of data, something of the utmost importance in the context of clustering. To overcome this, DLCC considers data depth based on depth-defined subsets of data. From this, local centers can be identified as well as clusters of varying shapes. Overall, DLCC is a flexible clustering approach that seems to overcome some limitations of traditional clustering methods, thereby enhancing data analysis capabilities across a wide range of application scenarios.
@@ I Introduction
The complexity of real-world data distributions has given rise to numerous clustering algorithms with different assumptions and objectives. Across those methodologies, many clustering approaches rely on identifying representative points (exemplars), such as centers, density peaks, or anchors, and forming clusters by grouping these exemplars and assigning remaining points accordingly. In the following, we review representative clustering methods under this unified viewpoint.
@@ I-A1 Center-based Clustering
Center-based clustering methods form clusters around a set of representative points defined as cluster centers, typically obtained by minimizing within-cluster dispersion under a prescribed distance measure. The classical k-means algorithm [9] exemplifies this paradigm, where exemplars are defined as arithmetic means of cluster members. While computationally efficient and conceptually simple, k-means is sensitive to outliers, biased toward spherical cluster shapes, and highly dependent on initialization, often leading to suboptimal local minima.
@@
Many variants of k-means have been proposed to address those limitations and refine the definition of exemplars. Partitioning around medoids (PAM; [10]) replaces means with medoids, improving robustness by restricting exemplars to sample points. Weighted k-means (W-k-means; [11]) and lasso weighted k-means (LW-k-means; [12]) provide a more flexible k-means approach by incorporating a weight vector for each variable, where LW-k-means enables feature selection in high-dimensional data by adding a lasso penalty, partially alleviating the spherical symmetry assumption.
Beyond hard assignments, fuzzy extensions of k-means reinterpret exemplars as prototypes that attract data points with varying membership strengths. Classical fuzzy C-means [13] and its variants define cluster centers as membership-weighted averages, thereby softening hard boundaries while retaining a center-based exemplar structure. Equilibrium k-means (EKM; [14]), a recently proposed method, addresses the limitations of fuzzy clustering under class imbalance by modifying the interaction between data points and exemplars through an attraction–repulsion mechanism, resulting in more stable and representative centroid estimation.
Despite these advances, center-based methods fundamentally rely on distance-driven exemplar definitions and a one-exemplar-per-cluster assumption, which makes them sensitive to metric choice and limits their ability to model complex and nonconvex data geometries.
I-A2 Density-based Clustering
@@
Density-based clustering assumes that clusters are regions of high density separated by low-density areas, making it well-suited for detecting arbitrarily shaped clusters. In this paradigm, exemplars are implicitly or explicitly defined as density maxima or points located within dense regions. A classical example is density-based spatial clustering of applications with noise (DBSCAN; [15]), which characterizes clusters as density-connected components, where each component is represented by a “core point”. While effective for arbitrary shapes, its reliance on a fixed threshold limits its ability to handle clusters with varying densities.
Mean shift clustering [16] further adopts a density-driven exemplar definition by iteratively shifting points toward local modes of an estimated density function. The resulting modes act as exemplars that summarize local structure without requiring a predefined number of clusters. However, the effectiveness of mean shift depends critically on hyperparameters for estimating density (e.g., the kernel bandwidth).
Density peak clustering (DPC; [17]) explicitly defines exemplars as density peaks that are well separated from higher-density points. A key limitation of it is that spurious (satellite) peaks may be promoted to exemplars, which degrades robustness. Recent work has therefore incorporated hierarchical ideas to improve exemplar grouping. In particular, fast main density peak clustering within relevant regions (R-MDPC; [18]) introduces the notion of relevant regions by first forming single-peak clusters led by density peaks and then merging density-connected single-peak clusters into larger regions, within which main peaks are identified. This enhancement better handles complex cluster shapes and reducing satellite peak interference. More broadly, such hierarchical schemes provides a generic mechanism for grouping representatives via linkage criteria. See also [19, 21, 20] for related clustering methods that adopt hierarchical schemes to merge over-fragmented components.
Density-based approaches still depend on density estimation and/or distance-based relations among representatives, which can be unstable under strong overlap between clusters or in high-dimensional spaces.
I-A3 Graph-based Clustering
Graph-based clustering forms clusters by partitioning a similarity graph, where vertices represent samples and edge weights encode pairwise affinities. Spectral clustering (SC; [22, 23]) solves a relaxed graph-cut problem via Laplacian eigen-decomposition and typically applies a discretization step (e.g., k-means) to obtain hard labels.
To avoid potential information loss from post-processing discretization and to couple graph construction with label learning, constrained Laplacian rank (CLR; [24]) approaches directly produce discrete labels from an optimally structured graph. However, both SC and CLR-style formulations can be computationally prohibitive on large-scale data due to full-graph construction.
Exemplars in graph-based clustering are used not only to summarize data structure but also to reduce the computational burden of graph construction and partitioning. This has motivated anchor-based graph clustering, which selects $m\!\ll\!n$ representative anchors and builds a sparse bipartite graph to approximate the original affinity graph [25, 26, 27]. Although anchors improve scalability, the resulting graph can remain sensitive to the choice of distance metric, graph-construction hyperparameters, and the anchor selection strategy.
@@ I-B Motivation
Viewed through an exemplar-based lens, many clustering methods succeed or fail primarily due to how representativeness is defined. Most existing approaches tie exemplars to distance- or density-driven criteria (and their associated hyperparameters). Such criteria can work well when a single notion of “closeness” or “density” is appropriate, but may yield misleading exemplars when the chosen metric or density scale is mismatched to the data (e.g., Euclidean distance under highly skewed distributions, or DBSCAN-style density reachability in high-dimensional spaces).
On the one hand, center-based methods typically define a single exemplar per cluster through distance minimization, an assumption that can be overly restrictive for irregular or nonconvex structures. On the other hand, density-based methods often induce multiple exemplars (e.g., core points or density peaks), but the resulting exemplar quality depends on reliable density definition.
@@
These limitations motivate an exemplar definition that remains consistent across dimensionalities and geometries, while allowing the number of exemplars to adapt to local structure. In response, we propose the depth-based local center clustering (DLCC) algorithm. DLCC addresses this by leveraging statistical data depth, a nonparametric tool for multivariate analysis. Tukey [28] pioneered the idea of center-outward ordering in multivariate data, with the introduction of the first depth definition, half-space depth. Data depth enables the center-outward ordering of data, making it a potent tool for locating a center for any dataset. When considering subsets of the data, data depth with respect to (w.r.t.) each subset aids in defining multiple exemplars. Naturally, as the subset size grows, the number of exemplars diminishes, and vice versa.
@@
To inherit the simplicity of center-based clustering and the shape adaptivity of density-based approaches, DLCC combines depth-defined local exemplars with two grouping strategies for exemplar aggregation. Here, the term “depth-based” encompasses two aspects: (1) data depth is used to construct a similarity matrix instead of relying on distance, adaptively capturing the geometric characteristics of each point’s location; (2) data depth measures the centrality of points within each point’s neighborhood, helping to identify exemplars.
I-C Preliminaries about depth and related work in clustering
Serfling and Zuo [29] formulated several desirable properties of depth functions: affine invariance, maximality at center, monotonicity on rays, and vanishing at infinity. A deepest point $\mathbf{m}=\text{argmax}_{\mathbf{z}\in\mathbb{R}^{d}}D(\mathbf{x}\mid F)$ is usually referred to as a depth median. In applications, where the inherent distribution $F$ is usually unknown, depth is calculated w.r.t. the empirical distribution $F_{n}$ of $\mathbf{X}$, that is $D(\mathbf{z}\mid F_{n})$ is the depth of $\mathbf{z}$ w.r.t. the sample $\mathbf{X}$. Hereinafter, we will denote the latter as $D(\mathbf{z}\mid\mathbf{X})$ to emphasize the role of the sample on calculating the depth of the point $\mathbf{z}$. For a broader overview of various depth notions, see Mosler and Mozharovskyi [30].
with fixed depth median versions
where $\mathbf{z}$ denotes a point in $\mathbb{R}^{d}$, $\bar{\mathbf{X}}$ is the mean vector, $\mathbf{x}_{i}$ is an observation in $\mathbf{X}$, while $\hat{\mathbf{\Sigma}}$ is the estimated covariance matrix of dataset $\mathbf{X}$. Unlike DBSCAN, which defines neighborhoods via a fixed Euclidean radius, DBCA uses (2) to induce elliptical, affine-invariant neighborhoods. Subsequent extensions further developed this depth-based framework by refining robustness of covariance matrix estimation and rules of core points [33, 34].
Another recent development in depth-based clustering was proposed by Francisci et al. [35], who build upon the concept of local depth introduced by Agostinelli and Romanazzi [36]. Their work generalizes local depth to lens depth and $\beta$-skeleton depth, and exploits depth as a surrogate for density estimation to develop a mean-shift like clustering algorithm. Beyond density-based paradigms, data depth has also been explored for identifying representative centroids in center-based clustering, particularly for k-means initialization across diverse datasets [37, 38].
So far, the notion of depth serves an auxiliary role to the more traditional clustering methods and is not employed directly in examining the multimodality of data. In contrast, DLCC places data depth at the core of both exemplar identification and similarity construction, thereby using depth directly to examine local and global cluster structure.
The DLCC framework itself is not restricted to a specific depth function, and in principle can be instantiated with a broad class of depth notions. In this work, we adopt spatial depth (SD; [39]) as a representative choice, as it offers a favorable trade-off between theoretical interpretability and computational simplicity. Although SD is invariant only under similarity transformations, its straightforward computation makes it well suited for high-dimensional applications. The sample version of SD is given by
where $n$ is the total number of observations.
DLCC requires a similarity matrix to determine the neighbors of each observation. Traditional distance measures, such as Euclidean distance, are typically based on assumptions about the underlying cluster shapes (e.g., spherical or linear structures), which may not hold in more complex scenarios. In contrast, data depth offers a nonparametric, robust, and adaptive alternative. As the depth (3) provide a center-outward ordering and their value ranges from 0 to $1$, the depth values of other points can be viewed as a measure of similarity to the depth median. Thus, similarity matrices can be constructed based on data depth. For this, however, one requires an approach capable of defining neighbors of any point $\mathbf{x}_{i}$ in $\mathbf{X}$, i.e., the observations in $\mathbf{X}$ that are most similar to $\mathbf{x}_{i}$. It should be clear that the usual notion of depth cannot be applied directly here, as that can only provide similarity measurements related to the global median.
We note that this new set is comprised of all the points belonging to the original sample $\mathbf{X}$ as well as of the reflection of all the points of $\mathbf{X}$ through $\mathbf{x}_{i}$. Crucially, $\mathbf{x}_{i}$ is a depth median of $\mathbf{X}_{Ri}$ as it is a center of symmetry for this set. It is also the case that $D(\mathbf{x}_{i}|\mathbf{X}_{Ri})=1$ and that $D(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ can be used to quantify the similarity of any point $\mathbf{x}_{j}$ ($j\neq i$) to the point of interest $\mathbf{x}_{i}$. Subsequently, by repeatedly computing the depth values for points in $\mathbf{X}$ w.r.t. $\mathbf{X}_{R1},\mathbf{X}_{R2},\ldots,\mathbf{X}_{Rn}$, a $n\times n$ non-mutual depth-based similarity matrix can be defined. We denote it as $\mathbb{S}$, where $\mathbb{S}_{i,j}=D(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ represents the entry in the $i$th row and $j$th column. As will be seen shortly, this can cause some computational challenges.
This approach to localizing the notion of depth was firstly introduced by Paindaveine and Van Bever [40], who proposed a form of local depth that differs from the one of Agostinelli and Romanazzi [36]. In particular, their construction was shown to yield non-spherical depth neighborhoods in some contexts. The same authors [41] used this concept and introduced an affine-invariant k-nearest neighbor classification algorithm that adapts to the local geometry of the sample and was shown to be consistent.
I-D Main Contributions and Organization of the Paper
@@
  • Efficient Depth-Based Similarity Construction: We improve the computational efficiency of constructing depth-based similarity matrices under spatial depth by developing a matrix-based implementation.
  • DLCC Clustering Framework: We propose DLCC, a depth-based clustering framework that uses data depth to measure similarities, define robust neighborhoods, and extract exemplars (“local centers”). DLCC then forms clusters by grouping these local centers and propagating labels to all observations.
  • A Complete Procedure and Empirical Evaluation: We provide a practical end-to-end procedure for DLCC, including filtering, grouping local centers and cluster formation mechanisms, together with guidance on strategy/parameter selection. We also provide extensive experiments on synthetic and real-world data sets.
The remainder of the paper is organized as follows: Section II introduces the depth-based similarity matrix. Section III provides a detailed description of the DLCC framework. In Section IV, we present details about parameter selections. Section V showcases experimental results of DLCC using both synthetic and real-world datasets. Lastly, Section VI concludes the paper.
@@
DatasetDepth-basedsimilarity matrixSection IINeighborhoods &Local centersSection III Filtered centersSection III-AGroups of filtered centersSection III-A &\& Section IV-BTemporaryclustersSection III-BUnlabelled pointsSection III-BFinal clustersclassification
@@
$D_{\mathrm{SD}}(\mathbf{z}|\mathbf{X})$ Spatial depth of $\mathbf{z}$ w.r.t. $\mathbf{X}$;
@@
$\mathbb{S}$ Depth-based similarity matrix, with $\mathbb{S}_{i,j}$ its $ij$th entry;
@@
$c$, $c$, $\mathbf{C}$ A local center, a certain set of local centers, and the set of all filtered centers, respectively;
$\mathbb{M}$ Local center neighborhoods’ similarity matrix, with $\mathbb{M}_{i,j}$ its $ij$th entry;
@@
$K$ and $k$ Number of clusters and cluster index;